Saturday, 22 August

Other Events

\underline{18:00 Urania} Movie

Errors aren’t Forever / Law of Least Effort / The new Shepherd’s Lamp

\underline{19:30 Urania} Jacobus van Lint, Eindhoven, Netherlands

The Mathematics of the CD–Player

\underline{20:00 Urania} VideoMath Festival

The “VideoMath Reel”, the 24 best contributions to the VideoMath competition

ICM’98 – Program  1